OpenStudy (anonymous):

A bus travels 250 km south along a straight path with an average velocity of 81 km/h to the south. The bus stops for 26 minutes then it travels 230 km south with an average velocity of 78 km/h. What is the average velocity for the total trip?

OpenStudy (anonymous):

firstly find the time taken by bus to cover teh 250 km and then find the time to cover the 230 km and then add the total time also add 26 minutes . after that add both distances 250 and 230 and divide by total time . you will get your average velociity.

OpenStudy (anonymous):

250/81 we get time for first travel then we 230/78 now we get the time for second travels then add these time and add bos stop time. After that add the distance then for average velocity total distance divide by total time. you will get average velocity.

OpenStudy (anonymous):

250/81 = x 230/ 78 =y then add total time : x + y + 26/60 ( convert 26 min into hours) now add total distance: 250 + 230 = 480 km now average velocity is : 480 / total time

OpenStudy (anonymous):

Thank you!!! ^.^

OpenStudy (anonymous):

welcome dear.

OpenStudy (anonymous):

welcome..

OpenStudy (anonymous):

This isn't in km/h is it?

OpenStudy (anonymous):

the answer will be in km/ h.

OpenStudy (anonymous):

480 km/h wasn't right

OpenStudy (anonymous):

divide 480 by total time than you will get your answer.

OpenStudy (anonymous):

Thanks!! I got it! :)

OpenStudy (anonymous):

welcome dear